# V2X Cooperative Perception for Autonomous Driving: Recent Advances and Challenges

### v3

@@ VII Lessons Learned
CP technology plays a pivotal role in overcoming the limitations inherent in single autonomous vehicles, addressing issues like restricted perception range and obscured areas. However, the unique characteristics of autonomous driving and the evolution of communication technology bring forth a multitude of challenges for CP. Drawing from recent research and comprehensive analysis of cooperative autonomous driving, we have gleaned valuable insights and identified key open challenges that will significantly influence the future research trajectory in this field. In this section, we will delve into the foremost among these insights, providing an in-depth examination of the hurdles that V2X CP must overcome to fully realize its anticipated potential.
VII-A Open Challenges
VII-A1 Collaboration Trigger
Much of the existing literature predominantly focuses on delineating cooperative behaviours among agents within specific scenarios, often neglecting the essential aspect of empowering the ego CAVs to discern opportune moments and suitable contexts for collaboration. To facilitate this, CAVs require a collaboration trigger mechanism capable of ascertaining the potential benefits of solo travel versus cooperative interaction with other agents at a given time step and within the prevailing driving conditions. The construction of effective collaboration triggers requires exploring pertinent decision factors, including but not limited to weather conditions, road terrain characteristics, and the state of onboard sensors. Furthermore, it is imperative to consider the temporal aspect of cooperation, ensuring that its duration aligns with the preservation of string stability among multiple agents. A compelling avenue for further research involves investigating whether the advantages of cooperation should be anticipated and communicated to CAVs in advance and whether this proactive knowledge dissemination would influence the overall performance of cooperative endeavours. In light of these research gaps, future endeavours are encouraged to tackle the complexities inherent in cooperative decision-making processes.
VII-A2 Cooperative Perception Under Constraints of Realistic Communication
Existing methods for CP have traditionally focused on optimizing the perception capabilities, often disregarding or simplifying the critical constraints imposed by communication factors in the real world. Yet, the real-world communication constraints can heavily degrade the CP performance. Thus, such constraints have to be mimicked realistically. For instance, the limited communication bandwidth emerges as a pivotal constraint, significantly impacting real-time collaboration by increasing transmission time in practical scenarios or reducing the number of bits that can be shared over wireless link. Furthermore, shared data is susceptible to interference from other agents during transmission, exacerbated by obstacles between transmitters and receivers, causing multi-path effects and fluctuation of signal power [285]. Selecting agents with sub-optimal communication interference and poor-quality wireless propagation channels can result in a diminished Signal-to-Interference plus Noise Ratio, reduced data rates, and elevated Bit Error Rates. Consequently, the information transmitted for CP may be discarded or incorrectly decoded. In light of these challenges, it becomes imperative to devise CP technologies that holistically integrate realistic communication constraints modelled by proper networking simulators, detailed-oriented cross-layer communication system models and more reliable wireless signal propagation models [286, 287], alongside the goal of maximizing perception capabilities in practical applications.
VII-A3 Malicious and Selfish Behavior
Reliable communication is of utmost importance to foster effective collaborative driving. Nevertheless, it remains plausible that some agents may occasionally engage in self-serving behaviour, collaborating solely to reduce their costs while causing detriment to others. These self-interested individuals, commonly referred to as “free-riders”, exploit the information shared by other agents to reap the benefits of coordinated efforts without contributing their insights and observations to the collective knowledge pool. Moreover, the distributed deployment of multiple agents diminishes defences against cyber-attacks, rendering them susceptible to malicious activities initiated by untrustworthy users [288]. A malicious collaborator may intentionally transmit erroneous environmental data to neighbouring agents, purposefully dispatch excessive computational tasks, and exhaust valuable computing resources and network bandwidth. The actions of malicious and self-interested agents pose a significant threat to the integrity of the vehicle communication network. The resulting degradation in network services not only undermines communication reliability but also jeopardizes the safety of autonomous vehicles. To fortify the security of the system network, it is imperative for future research to place a heightened emphasis on trust-related concerns among agents. This entails the identification and mitigation of malicious and self-serving behaviours by incorporating them into CP schemes, ideally through the application of suitable methodologies like game theory [289].
VII-A4 Generalizability in Real Scenarios
Constructing real-world scenes and collecting data for researching CP techniques presents substantial challenges. Consequently, many CP methods rely on simulated datasets for development. Although a few recent real-world datasets have become available [273, 275, 274], they are often limited in the number of participating agents, failing to capture the complexities of real-world collaboration scenarios. Therefore, CP frameworks trained on datasets generated through simulation tools may lack optimal generalizability in realistic collaboration settings. Regrettably, the outlook for collecting and annotating perception data in genuine V2X scenarios remains uncertain in the near term. As a viable alternative, it is essential to explore the development of feasible domain adaptation methods [290]. These methods would enable the transfer of CP models between authentic and simulated domains, ultimately facilitating the practical application of V2X CP in various scenarios.
VII-A5 Challenging Scenes and Corner Cases
@@
Ensuring the safety of V2X autonomous driving hinges on the ability to attain robust and precise perception results across diverse and complex scenarios. While several large-scale CP datasets have emerged in recent years, they predominantly concentrate on typical traffic scenarios, often overlooking challenging scenes, notably corner cases. Corner cases encompass atypical and extreme situations within autonomous driving scenarios, marked by uncommon conditions, hardware impediments, and irregular obstacles such as overturned accident vehicles or uniquely shaped vehicles. Due to their scarcity, these corner cases are frequently omitted from perception model training for autonomous driving. Consequently, these models struggle to accurately identify objects, resulting in instances of either missed or erroneous detections, ultimately contributing to severe traffic accidents. To address this pressing issue, there is an immediate need to introduce CP datasets that encompass intricate scenarios and corner cases, bolstering the development of a more resilient CP system.
VII-B Future Direction
VII-B1 Hybrid Collaboration Scheme
Efficient collaboration in multi-agent systems is crucial for CDAs. Early collaboration preserves vital information but requires high-speed communication due to raw data size. Late collaboration with object-level data reduces data size but may degrade performance by discarding valuable information during object proposal merging. Intermediate collaboration, focusing on feature-level data transmission and integration, balances accuracy and bandwidth consumption. However, persistent bandwidth saturation or underutilization challenges exist.
One promising solution involves designing an efficient collaboration trigger. This trigger determines when to cooperate with other agents and selects the most suitable collaborative scheme. Driven by intelligent algorithms and adaptive decision-making, it optimizes bandwidth and resource utilization. Additionally, the hybrid collaboration technique, integrating three fusion schemes, enables agents to partition input frames into raw, feature, and object data segments, facilitating exchange with collaborators. This approach capitalizes on diverse fusion techniques’ strengths and adapts to variable network conditions and environmental contexts. Combining intelligent collaboration triggers with the hybrid collaboration technique holds transformative potential, enhancing multi-agent system efficiency and performance in complex, dynamic scenarios with improved precision and resource management.
VII-B2 Multi-modal Data Sharing
LiDAR sensors offer precise geometric data, while cameras provide rich semantic information. Therefore, a multi-modal perception system could theoretically improve performance by leveraging complementary sensor data. In cooperative autonomous driving, the focus has been mainly on collaboration among agents equipped with identical sensors. However, using different sensor types across various agents can reduce costs while maintaining desirable performance. Integrating multi-modal data from multiple agents presents systematic challenges due to dynamic geometric relationships between LiDAR and camera inputs. Research into CP involving the sharing of multi-modal data is valuable. It is important to note that not all agents possess a suite of multi-modal sensors. Hence, careful consideration of sensor modality is essential when selecting collaborative agents. For instance, prioritizing agents capable of providing high-resolution data for collaboration is advisable during adverse weather conditions.
VII-B3 Integrated Sensing and Communication for Cooperative Perception
@@
ISAC technology can help overcome challenges associated with existing cooperative agent selection methods. It enables dynamic bandwidth allocation and adaptable communication strategies, mitigating limitations imposed by bandwidth constraints, interference, and multi-path effects. This adaptive optimization process ensures efficient and reliable data transmission to its intended destination. Furthermore, ISAC can incorporate error correction and data integrity mechanisms to minimize the risk of data loss or incorrect decoding during transmission. Therefore, ISAC aligns with the need for comprehensive agent selection schemes considering both sensing and communication aspects. It aims to optimize perceptual capabilities while addressing practical challenges related to limited communication resources and potential interference. Ultimately, this advancement enhances the effectiveness of collaborative perception in complex environments.
VII-B4 Responsible AI for Cooperative Perception
Cooperative autonomous driving relies heavily on Artificial Intelligence (AI), which excels in mastering increasingly complex computational tasks. However, AI’s use can also raise concerns, including privacy breaches, bias, and opacity. Balancing AI’s promises with its risks necessitates the adoption of responsible AI principles. These principles, such as fairness, transparency, and privacy, provide guidelines for implementing AI models. While there is a growing focus on responsible AI frameworks, their practical application and impact in V2X CP design have received limited attention. The translation of high-level principles into concrete AI design practices within CP remains unclear.
Responsible AI initiatives can offer valuable insights for creating realistic simulation environments that closely resemble real-world scenarios, benefiting the training and validation of V2X CP frameworks. Additionally, these initiatives can drive the development of domain adaptation techniques, enhancing knowledge transfer between simulated and real domains. This knowledge transfer improves the generalisation of the CP model. Moreover, responsible AI efforts can stimulate the advancement of deep learning methods tailored for generating CP data featuring intricate scenarios and corner cases. Generative models and ethically guided data augmentation techniques are pivotal for constructing comprehensive datasets. This enriched dataset forms the basis for robust training for the V2X CP system, equipping them to navigate and respond effectively to challenging real-world situations.
VII-B5 Privacy-preserved Cooperative Perception
To establish sustainable CP, ensuring the privacy of data and messages exchanged via V2X communications is of paramount importance [292]. Traditional CP systems often rely on centralized servers like RSUs or master vehicles for data fusion algorithms. However, data sharing over V2X communications poses potential risks, as sensitive user information may become exposed to external threats or unauthorized access by servers. FL offers a viable solution to enhance CP intelligence while preserving privacy. FL facilitates intelligent CP by enabling fusion algorithms to run on servers without requiring access to raw data or information sharing. It employs a distributed training mechanism in which collaborating vehicles train localized CP models using their data and then share only the trained model with the server for global aggregation [293]. This approach ensures secure distributed learning, safeguarding the shared model from external adversaries. By prioritizing privacy preservation, FL encourages more CAVs to participate in data perception, thus enhancing the overall robustness of CP systems.
VII-B6 Collaborative AI Integration for Enhanced CP and V2X Communication
@@
Consequently, further research is imperative to orchestrate effective collaboration and coordination among distinct AI systems operating within individual CAVs and AI systems operating within the V2X network infrastructure. These AI systems must align their objectives and strategies to achieve efficient CP, particularly in dynamic and complex traffic scenarios. Such endeavours are essential to harness the full potential of AI-driven CP for enhancing road safety and traffic efficiency.
@@
VII-B7 Migration of Advanced Perception and Autonomous Driving Framework from Single Vehicle to Cooperative Vehicles
@@
As the development of transformer model [295] in both computer vision and large language model areas, several advanced technologies have been proposed for single-vehicle perception and autonomous driving, e.g. end-to-end perception [296], occupancy grid prediction [297], referring expression-based perception [298], self-supervised world-models for autonomous driving [299, 300], to boost the downstream tasks by using the features extracted from understanding and predicting/reconstruction of the environment, and end-to-end autonomous driving [301], etc. All of aforementioned technologies have great potentials to be extended into cooperative vehicles setting as well, with the expectation for enhancing the capability of autonomous driving. The most recent example is end-to-end autonomous driving under V2X cooperation [302], in which the end-to-end autonomous driving framework is integrated with V2X cooperation to achieve significant performance improvement on planning and other intermediate outputs.
VIII Conclusions
In this paper, we thoroughly examined V2X-based CP and its recent technological progress. Our first focus was on the advancements in V2X communication technologies that support the CP in CAVs. We specifically reviewed recent developments in edge computing, blockchain, digital twin, and 6G that provide low latency, high data rate, and reliable and secure information transmission for CP. Then, we presented a comprehensive CP framework that encapsulates these systems’ fundamental components and core aspects. By adopting a new taxonomic classification for mainstream architectures and conducting in-depth reviews of the literature, we dissected the intricate cooperative challenges that lie at the heart of CP. Furthermore, we highlighted the availability of open-source databases and simulators, which are invaluable tools for evaluating and validating the performance of CP models. These resources facilitate rigorous testing and refinement of CP algorithms in a controlled environment. Moreover, we emphasized the importance of incorporating ethical principles such as fairness, transparency, and privacy into designing and implementing AI models for CP. These principles ensure the ethical deployment of AI in cooperative driving and contribute to the creation of highly realistic simulation environments, domain adaptation methodologies, and enriched datasets that can empower CP models to navigate the complexities of real-world scenarios effectively. In summary, CP is a rapidly evolving area that holds great promise. By combining state-of-the-art V2X communication technology, multi-modal sensing technology, and AI technology for data processing, reliable and efficient CP can be achieved, thus creating safer, more efficient, and more capable autonomous driving and intelligent transportation systems.

### v4

@@ VIII Lessons Learned
CP technology plays a pivotal role in overcoming the limitations inherent in single autonomous vehicles, addressing issues like restricted perception range and obscured areas. However, the unique characteristics of autonomous driving and the evolution of communication technology bring forth a multitude of challenges for CP. Drawing from recent research and comprehensive analysis of cooperative autonomous driving, we have gleaned valuable insights and identified open challenges and future directions that will significantly influence the research trajectory in this field.
VIII-A Challenges and Potential Solutions
VIII-A1 Collaboration Trigger
A prevailing challenge within CP is the ability of each CAV to autonomously identify optimal opportunities for collaboration. Current research frequently emphasizes predefined cooperative behaviors, often neglecting the necessity for ego CAVs to dynamically evaluate when collaboration is advantageous. This process necessitates the development of a mechanism referred to as a “collaboration trigger,” which assesses the trade-offs between independent actions and collaborative efforts, informed by real-time conditions such as weather, road attributes, and sensor performance. Furthermore, ensuring stable and temporally consistent cooperation among multiple CAVs presents significant challenges, particularly in the context of dynamic driving scenarios.
To address these challenges, adaptive, context-aware systems are needed to determine when collaboration is optimal. A dynamic decision model allows each ego CAV to assess real-time factors like sensor performance, road conditions, and traffic, ensuring stable and efficient cooperation. Key metrics, such as perception accuracy and energy efficiency, can guide these decisions, with a fusion algorithm dynamically weighing factors like weather and traffic forecasts as conditions change. Effective implementation requires a cloud or edge-computing platform to aggregate real-time and historical data, enabling CAVs to make informed, anticipatory collaboration choices.
VIII-A2 Real-world Communication Constraints
Existing CP methods have traditionally focused on optimizing perception performance, often oversimplifying the critical challenges posed by real-world communication constraints such as limited bandwidth, unstable channels, network congestion, and data delays. These factors compromise the data received by the ego vehicle, and unlike conventional applications, retransmitting data in CP can render it unusable due to outdated information [316]. This results in missing data, asynchronous information from other agents, and, ultimately, inaccurate data fusion within the CP system.
A promising approach is the integration of cross-layer models to assess how communication constraints affect CP algorithms [317, 318]. For example, combining network-layer routing with physical-layer channel modeling enables CP systems to anticipate the effects of channel fading or congestion on data transmission, facilitating adaptive adjustments in CP processes. Additionally, predictive perception models can enhance CP resilience by using historical data and probabilistic modeling to forecast surrounding vehicle behavior, maintaining situational awareness even during data loss. This reduces dependency on continuous communication, ensuring reliable CP performance in real-world conditions.
VIII-A3 Malicious and Selfish Behavior
Trustworthy communication is essential for fostering effective and safe CP. However, some road users may act in self-serving ways, collaborating only to reduce their own costs while potentially harming others. These self-interested users, commonly known as “free-riders,” exploit shared information to benefit from coordinated efforts without contributing their own insights to the collective knowledge pool. Furthermore, the distributed deployment of multiple agents lowers defenses against cyber-attacks, increasing vulnerability to malicious activities by untrustworthy users [319]. Malicious collaborators may deliberately transmit inaccurate environmental data, compromising the integrity of CP by degrading data quality and reliability. This degradation not only disrupts CP effectiveness but also heightens risks to the safety of CAVs that rely on accurate, shared situational awareness.
To tackle the challenges posed by unreliable collaborators in CP, several strategies can enhance data reliability and safety. Implementing trust scoring and reputation systems allows CAVs to prioritize data from agents with a history of reliability. Machine learning-based anomaly detection can also identify unusual patterns that might indicate malicious behavior. Cross-verifying data from multiple sources and utilizing multi-modal sensor fusion further improves the accuracy of CP, enabling CAVs to detect and disregard inconsistent or unreliable inputs. Blockchain technology can introduce a decentralized layer of trust by immutably logging data interactions, increasing agent accountability. Lastly, dynamic data prioritization ensures that critical information, such as immediate obstacles, is prioritized, reducing the potential impact of harmful data. Together, these approaches strengthen the resilience of cooperative perception, supporting safer and more reliable collaborative driving.
VIII-A4 Generalizability in Real Scenarios
Constructing real-world scenes and collecting data for CP research presents substantial challenges, leading many CP methods to rely on simulated datasets for development. While some recent real-world datasets have emerged [292, 293, 294, 295, 296, 297], they typically involve limited agent participation, which fails to capture the full complexity of real-world collaborative scenarios. As a result, CP frameworks trained primarily on simulated datasets may struggle to generalize effectively in realistic collaboration settings.
To address this gap, domain adaptation and synthetic data enhancement techniques offer promising solutions that can improve the applicability of CP models to real-world environments. Domain adaptation models, for instance, align features between simulated and real-world datasets through adversarial training and transfer learning, reducing distributional discrepancies and enabling CP models to perform more reliably in diverse, real-world settings [320].
In addition to domain adaptation, synthetic data enhancement techniques improve the realism of simulated datasets. One approach, described in [321], synthesizes training data using Augmented Reality, then refines it through a Generative Adversarial Network (GAN) to create robust photo-realistic images under various weather and lighting conditions. Other generative models, such as diffusion models, also play a vital role by transforming simulated data into realistic visuals. These models adjust textures, lighting, and environmental effects, effectively bridging the gap between synthetic and real-world visuals. By enhancing dataset realism, these techniques support more robust CP model training and improve generalization across varied real-world scenarios.
VIII-A5 Challenging Scenes and Corner Cases
@@
While several large-scale CP datasets have been developed in recent years, many focus primarily on typical traffic situations and often neglect challenging scenarios, particularly corner cases. Corner cases are uncommon and extreme situations in autonomous driving, characterized by unusual conditions, hardware limitations, or irregular obstacles, such as overturned vehicles or unusually shaped objects. These rare scenarios are frequently excluded from perception model training, which hinders the model’s ability to identify objects accurately and can lead to missed or erroneous detections. Such inaccuracies in perception increase the risk of severe traffic accidents, highlighting the urgent need for CP datasets that incorporate both intricate scenarios and corner cases to support the development of more robust perception systems.
A combined approach of targeted data collection and synthetic data generation is essential to address the gap in corner cases within CP datasets. Simulation environments can be utilized to create synthetic data that replicates corner cases, including overturned vehicles, unusual object shapes, and scenarios with sensor occlusions. This method allows for the generation of a wide variety of challenging conditions without relying on rare real-life occurrences. Additionally, a model-driven data collection approach can be implemented, where trained CP models actively monitor their predictions in real-time. They can identify and flag low-confidence detections or false positives/negatives as potential corner cases. These flagged instances can then be collected and prioritized for further training. By creating datasets that encompass both typical and rare autonomous driving conditions, CP systems can be better trained and validated, ultimately improving performance and reliability in real-world situations.
VIII-B Future Directions
VIII-B1 Integrated Sensing and Communication for CP
@@
VIII-B2 Responsible AI for CP
Cooperative autonomous driving relies heavily on Artificial Intelligence (AI), which excels in mastering increasingly complex computational tasks. Responsible AI initiatives can offer valuable insights for creating realistic simulation environments that closely resemble real-world scenarios, benefiting the training and validation of V2X CP frameworks. Additionally, these initiatives can drive the development of domain adaptation techniques, enhancing knowledge transfer between simulated and real domains. This knowledge transfer improves the generalisation of the CP model. Moreover, responsible AI efforts can stimulate the advancement of deep learning methods tailored for generating CP data featuring intricate scenarios and corner cases. Generative models and ethically guided data augmentation techniques are pivotal for constructing comprehensive datasets. This enriched dataset forms the basis for robust training for the V2X CP system, equipping them to navigate and respond effectively to challenging real-world situations.
VIII-B3 Privacy-preserved CP
To establish sustainable CP, ensuring the privacy of data and messages exchanged via V2X communications is of paramount importance [323]. FL offers a viable solution to enhance CP intelligence while preserving privacy. FL facilitates intelligent CP by enabling fusion algorithms to run on servers without requiring access to raw data or information sharing. It employs a distributed training mechanism in which collaborating vehicles train localized CP models using their data and then share only the trained model with the server for global aggregation [324]. This approach ensures secure distributed learning, safeguarding the shared model from external adversaries. By prioritizing privacy preservation, FL encourages more CAVs to participate in data perception, thus enhancing the overall robustness of CP systems.
VIII-B4 Collaborative AI Integration for Enhanced CP and V2X Communication
@@
Consequently, further research is imperative to orchestrate effective collaboration and coordination among distinct AI systems operating within individual CAVs and AI systems operating within the V2X network infrastructure. These AI systems must align their objectives and strategies to achieve efficient CP, particularly in dynamic and complex traffic scenarios. Such endeavours are essential to harnessing the full potential of AI-driven CP to enhance road safety and traffic efficiency.
@@
VIII-B5 Migration of Advanced Perception Framework from Single Vehicle to Cooperative Vehicles
@@
As the development of transformer model [326] in both computer vision and large language model areas, several advanced technologies have been proposed for single-vehicle perception and autonomous driving, e.g. end-to-end perception [327, 45], 3D occupancy grid prediction [328, 329], referring expression-based perception [330] with multi-modal large language model [331], self-supervised world-models for autonomous driving to boost the downstream tasks by using the features extracted from understanding and predicting/reconstruction of the environment [332, 333], and end-to-end autonomous driving [334, 335], etc. All of the aforementioned technologies have great potential to be extended into cooperative vehicle settings as well, with the expectation of enhancing the capability of autonomous driving. The most recent examples are [336, 337] and [338, 339, 340], in which the end-to-end autonomous driving framework and 3D occupancy grid prediction function are integrated with V2X cooperation, respectively, to achieve significant performance improvement on planning and intermediate perception outputs.
IX Conclusions
In this paper, we presented a comprehensive framework for V2X CP, addressing its foundational components and key aspects. Through a novel taxonomic classification of mainstream architectures and an in-depth literature review, we analyzed the multifaceted challenges inherent to effective CP. We examined how V2X communication limitations, such as latency and data loss, can impact CP performance, emphasizing the need for resilient communication strategies to maintain reliable perception sharing. We also highlighted the value of open-source databases and simulators, which serve as essential tools for evaluating and refining CP models in controlled environments. These resources enable rigorous testing and iterative optimization of CP algorithms, promoting advancements in model accuracy and robustness. Furthermore, we underscored the importance of embedding ethical principles, such as fairness, transparency, and privacy, into the development and deployment of AI models for CP. Incorporating these principles fosters ethical AI usage in cooperative driving and supports the creation of realistic simulation environments, domain adaptation techniques, and diverse datasets, which are crucial for equipping CP models to handle the complexities of real-world scenarios. In conclusion, CP is a rapidly evolving field with significant potential. By combining advanced V2X communication technologies, multimodal sensing, and AI-driven data processing, CP offers a pathway toward reliable, efficient, and ethically sound solutions for autonomous driving and intelligent transportation systems, ultimately enabling safer, more effective, and highly capable autonomous mobility.